How would you solve these two problems by doing subsitution 2x-3y=9 and 5x+2y=-6

Answer by Maths68(1474)

2x-3y=9........(1)
5x+2y=-6.......(2)
Take (1)
2x-3y=9
2x=9+3y
2x/2=(9+3y)/2
x=(9+3y)/2
Put the value of x in (2)
5((9+3y)/2)+2y=-6
Multiply by 2 both sides
2*(5((9+3y)/2)+2*2y=-6*2
5(9+3y)+4y=-12
45+15y+4y=-12
19y=-12-45
19y=-57
19y/19=-57/19
y=-3
Put the value of y in (1)
2x-3y=9
2x-3(-3)=9
2x+9=9
2x=9-9
2x=0
2x/2=0/2
x=0

x,y=0,-3

Check
=====
Put the value of x and y in (1)
2x-3y=9
2(0)-3(-3)=9
9=9
Put the value of x and y in (2)
5x+2y=-6
5(0)+2(-3)=-6
-6=-6